### v1

# Optimal Control of Discrete-Time Nonlinear Systems

Abstract
This paper focuses on optimal control problem for a class of discrete-time nonlinear systems. In practical applications, computation time is a crucial consideration in solving optimal control problem, especially when meeting the real-time requirements of control systems. To address this challenge, this study proposes a novel framework based on the optimal control method. Firstly, the original optimal control problem is transformed into an equivalent optimization problem, which is resolved using the Pontryagin’s maximum principle from the perspective of optimal control. Furthermore, explicit formulas for computing both the gradient and hessian matrix of the cost function are proposed. A numerical algorithm with rapid convergence and high efficiency is also provided.
Index Terms:
Nonlinear optimal control, maximum principle, forward and backward difference equations.
I Introduction
Optimal control problem (OCP) is fundamentally an optimization problem (OP) focused on determining the control inputs for a dynamical system to minimize or maximize a cost function over time. It can be argued that many natural processes are essentially driven by optimization, reflecting the universal principle of optimality [1]. In a world increasingly shaped by optimization-based control, the growing diversity of applications demonstrates its essential role and wide-ranging impact across various fields. For example, the optimization-based control strategies have enabled bipedal, quadrupedal, and humanoid robots to walk reliably in standard environments, leading to their first practical deployments, such as Boston Dynamics’ Spot robot [2]. These advancements, along with the state-of-the-art performance of optimal control in various applications, emphasize the increasing importance of theoretical research in this field.
Viewed from a static perspective, the nonlinear OCP can be solved using a variety of existing optimization algorithms [3]. Model predictive control (MPC) represents a practical implementation of this theory, effectively transforming the complex OCP into a series of smaller optimization problems that are solved in real-time [4]. In autonomous driving scenarios, MPC was used to optimize the angular velocity and steering wheel angles of automatic guided vehicles (AGV) for improving tracking accuracy [5]. In [6], the control strategy for AGV parallel parking was built around MPC, which ensured smooth transitions, stability, and effective handling of non-convex state constraints. To reduce the computational demands of real-time optimization, an event-triggered MPC approach was applied to study path tracking control in AGV [7]. Although there are many available optimization algorithms, the complexity of the optimization problem often increases with longer terminal time, leading to a greater computational burden.
Conversely, considering a dynamic perspective, the Pontryagin’s maximum principle is the key mathematical tool for solving nonlinear OCP, focusing on the solution of highly coupled nonlinear forward and backward difference equations (FBDEs), which reveal the essence of optimal control. In [8], the decentralized optimal control and stabilization problems for linearly interconnected systems with asymmetric information were effectively addressed by solving stochastic FBDEs. Unfortunately, analytical solutions to nonlinear FBDEs are nearly impossible to obtain, except in the case of linear systems. A well-known numerical method for solving nonlinear FBDEs is the method of successive approximation (MSA) [9]. In the field of machine learning, [10] proposed a self-healing neural network architecture that employed this method to automatically detect and fix the vulnerability issue, thereby enhancing the robustness of the neural network. However, the direct application of MSA, a gradient-based method, often results in a slow convergence rate during computation.
To overcome the limitations of the aforementioned methods, [11] introduced an approach that combines nonlinear optimization with the maximum principle to improve the efficiency of solving the OCP. This method primarily employs FBDEs for gradient computation, replacing the traditional gradient computation in optimization. In [12], the hybrid gradient-based method was applied to simultaneously optimize unmanned aerial vehicle trajectories and control signals, minimizing mission completion time and energy consumption while ensuring smooth and practical implementation. Similarly, the control gains of a controller based on a fully actuated system approach were optimized using the hybrid method to ensure high-precision trajectory tracking while satisfying angular and actuator constraints [13]. Besides, the approach has also been extended to address the OCP of switched nonlinear systems [14, 15]. It should be noted that both the selection of optimization algorithms and gradient computation are crucial for computational efficiency, with each playing an indispensable role. However, most existing studies primarily focus on gradient computation, placing relatively less emphasis on the design of optimization algorithms.
In this paper, the OCP for discrete-time nonlinear systems is considered. A novel solution framework is introduced to address this problem. Unlike previous research that relies on existing optimization algorithms [11, 12, 13, 14, 15], this work presents a novel optimization approach from the perspective of optimal control. While it extends the findings of [16] and [17], which redefine the optimization problem and propose a superlinear convergent iterative algorithm, this study specifically addresses dynamic problems, focusing on the challenges associated with effectively computing the gradient and hessian matrix. To tackle these challenges, a new approach based on the maximum principle is introduced for the first time, transforming the solution process into solving two sets of FBDEs. This provides a unified computational approach for developing higher-order optimal control algorithms. The contributions of this paper can be summarized as follows:
  • 1)
    The discrete-time nonlinear OCP is transformed into an OP. By constructing a novel OCP, a rapidly convergent iterative algorithm for solving the OP is presented, establishing a new framework for solving the OCP based on the principle of optimal control theory.
  • 2)
    A unified and explicit formula based on the maximum principle for computing the gradient and hessian matrix in OCP is proposed in this work, which reduces the computational burden. Furthermore, this method can be generalized to other second-order optimization algorithms.
The rest of this paper is organized as follows. In section II, the discrete-time nonlinear OCP is given. In section III, the control design and the iterative algorithm’s implementation and details are presented. Section IV concludes the work.
Notations: $\mathbb{R}^{n}$ denotes the n-dimensional Euclidean space; $A^{T}$ represents the transpose of matrix $A$; $\frac{\partial p}{\partial q}$ denotes partial derivative of function $p$ with respect to $q$. 0 denotes the zero vector.
II Problem Statement
Consider the following discrete-time nonlinear OCP
$\underset{u_{[0,N]}}{\text{min}}\ J_{N}=\sum_{k=0}^{N}\Phi(x_{k},u_{k}),$ (1)
[…]
$\hskip 24.70004ptx_{0}=\sigma,\ k=0,1,...,N,$
in which $\Phi\in\mathbb{R}$ is the running cost, $F\in\mathbb{R}^{n}$ is a given deterministic function, $x_{k}\in\mathbb{R}^{n}$ is the state variable, $u_{k}\in\mathbb{R}^{m}$ is the control variable, and $x_{0}$ represents the given initial state. It should be noted that the cost function (1) considered here is in the lagrange form, which encompasses the bolza form as a special case when $\Phi(x_{N},u_{N})$ is set to $\psi(x_{N})$ at $k=N$ [18].
Notice that the nonlinear dynamical system (2) can be reformulated as
$\begin{bmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{N}\end{bmatrix}=\begin{bmatrix}F(x_{0},u_{0})\\ F(F(x_{0},u_{0}),u_{1})\\ \vdots\\ F(F(x_{N-2},u_{N-2}),u_{N-1})\end{bmatrix},$ (3)
where the state $x_{k}$ can be regarded as a function of $,(x_{0},u_{0},u_{1},$$,u_{k-1})$. For convenience, denote $z=[u_{0}^{T},u_{1}^{T},...,u_{N}^{T}]^{T}$. Then the OCP (1)-(2) can be restated as an OP
[…]
Remark 1
The OCP (1)-(2) is equivalent to the OP (4). From the perspective of optimal control, the OCP (1)-(2) can be solved using the Pontryagin’s maximum principle, which involves solving highly coupled FBDEs. Additionally, from the perspective of nonlinear optimization, the OP (4) can be tackled using existing optimization method, such as gradient descent algorithm. In the following, the strengths of the aforementioned methods will be combined to present an efficient iterative numerical algorithm.
III Main Results
In this section, in order to solve the OP (4), the problem is reformulated as a new OCP from an optimal control perspective, and a rapidly convergent iteration algorithm is proposed. Subsequently, the explicit formulas for computing the gradient and hessian matrix of the cost function are proposed by solving the corresponding FBDEs.
III-A Optimal Control Design
In order to solve the optimal controller $u_{k}$ in (2), rather than directly solving the OCP (1)-(2), the focus is placed on its equivalent OP (4). In what follows, the OP (4) is reconsidered from a new perspective, namely that of optimal control.
Define the following new OCP:
$\underset{\upsilon_{[0,M-1]}}{\text{min}}J_{M}=\sum_{i=0}^{M-1}[J_{N}(x_{0},z_{i})+\frac{1}{2}\upsilon_{i}^{T}R\upsilon_{i}]+J_{N}(x_{0},z_{M}),$ (5)
[…]
$\hskip 28.30005ptz_{0}=\eta,\ i=0,1,...,M-1,$
where $z_{i}\in\mathbb{R}^{m(N+1)}$ and $\upsilon_{i}\in\mathbb{R}^{m(N+1)}$ are the state and control variables, respectively, $z_{0}$ is the initial state value, $R$ is the positive definite matrix, and the cost function (5) is based on that of OP (4).
In order to solve the OCP (5)-(6), the maximum principle is applied to derive the following equations:
$\gamma_{i}$ $=\gamma_{i+1}+\nabla J_{N}(x_{0},z_{i}),$ (7)
00 0 $=R\upsilon_{i}+\gamma_{i},$ (8)
where (7) is the costate equation with the terminal condition $\gamma_{M}=\nabla J_{N}(x_{0},z_{M})$, and (8) is the equilibrium condition. It is known that any solution to the OCP must satisfy the FBDEs. Therefore, the optimal control and state of the OCP (5)-(6) can be obtained by solving the FBDEs (6)-(8). Although FBDEs (6)-(8) can be numerically solved using the existing MSA, its convergence rate is relatively slow. To improve computational efficiency, an iterative numerical algorithm has been developed by combining the Taylor expansion with mathematical induction in [16, 17], as outlined below.
$z_{i+1}$ $=z_{i}-g_{i},$ (9)
$g_{i}$ $=(R+\nabla^{2}J_{N}(x_{0},z_{i}))^{-1}(\nabla J_{N}(x_{0},z_{i})+Rg_{i-1}),$ (10)
$g_{0}$ $=(R+\nabla^{2}J_{N}(x_{0},z_{i}))^{-1}\nabla J_{N}(x_{0},z_{i}),$ (11)
where $i=0,1,...,M-1$.
Remark 2
In the iterative algorithm (9)-(11), the update step $g_{i}$ requires inner loop iterations starting from $g_{0}$. It should be noted that the algorithm faces considerable challenges in handling dynamic problem, i.e., the OCP (1)-(2), due to the computational complexity involved in evaluating $\nabla J_{N}$ and $\nabla^{2}J_{N}$. The following subsections provide the explicit iterative solution algorithms. For details on the convergence analysis and the superlinear convergence rate of the algorithm (9)-(11), please refer to [16, 17].
III-B Computation of Gradient
From the OCP (1)-(2), it can be seen that the cost function $J_{N}$ is a complex function on the control variables $u_{k}$ to be optimized. To enhance computational efficiency, the gradient formula for the cost function $J_{N}$ is provided in this subsection.
To simplify presentation, a one-dimensional case for the control variables $u_{k}$ is considered, with a natural extension to higher dimensions. Define the following Hamiltonian
$H_{N}(x_{k},u_{k},\lambda_{k+1})=\Phi(x_{k},u_{k})+\lambda_{k+1}^{T}F(x_{k},u_{k}),$ (12)
where $k=0,1,...,N$, and $\lambda_{k}\in\mathbb{R}^{n}$ is the costate variable. Applying the Pontryagin’s maximum principle to the OCP (1)-(2), the main result is summarized in the following lemma.
Lemma 1
For the OP (4), the gradient of $J_{N}$ with respect to $z$ is given by
$\nabla J_{N}=\left[\frac{\partial H_{N}(x_{0},u_{0},\lambda_{1})}{\partial u_{0}},...,\frac{\partial H_{N}(x_{N},u_{N},\lambda_{N+1})}{\partial u_{N}}\right]^{T},$ (13)
where $x_{k}$ and $H_{N}$ are given by (2) and (12), and $\lambda_{k}$ is given by
$\lambda_{k}=\frac{\partial H_{N}(x_{k},u_{k},\lambda_{k+1})}{\partial x_{k}},$ (14)
with the terminal costate value $\lambda_{N+1}=\textbf{0}$.
Proof: The proof can be found in [11] and is omitted here.
The implementation steps of Lemma 1 are summarized as follows:
Step 1: For the given control $u_{k},k=0,1,...,N$, compute the forward difference equation (2) from $k=0$ to $k=N-1$ to obtain the state $x_{k},k=0,1,...,N$ using the given initial state value $x_{0}$.
Step 2: Compute the backward difference equation (14) from $k=N$ to $k=1$ to obtain the costate $\lambda_{k},k=1,...,N+1$ using the given terminal costate value $\lambda_{N+1}$.
Step 3: Compute the gradient of $J_{N}$ according to (13).
III-C Computation of Hessian Matrix
In this subsection, the second-order derivative of the cost function $J_{N}$ is computed based on Lemma 1, and the explicit formula for hessian matrix computation will be presented.
During the computation process, we solve for the elements of the hessian matrix line by line. For convenience, denote
$G(x_{k},u_{k},\lambda_{k+1})=\frac{\partial H_{N}(x_{k},u_{k},\lambda_{k+1})}{\partial x_{k}},$
$L(x_{k},u_{k},\lambda_{k+1})=\frac{\partial H_{N}(x_{k},u_{k},\lambda_{k+1})}{\partial u_{k}}.$
To compute the elements in the $i$th row of the hessian matrix, construct the following OCP:
$\underset{u_{[0,N]}}{\text{min}}\hskip 11.99998ptJ_{i}=L(x_{i},u_{i},\lambda_{i+1}),$ (15)
[…]
$\hskip 33.40001pt\lambda_{k}=G(x_{k},u_{k},\lambda_{k+1}),\ k=N,...,1,$ (17)
$\hskip 33.40001ptx_{0}=\alpha,\ \lambda_{N+1}=\textbf{0},$
where $i=0,1,...,N$, and the cost function (15) is subject to constraints imposed by the FBDEs (16)-(17).
Let $L(x_{i},u_{i},\lambda_{i+1})=\sum_{k=0}^{N}L(x_{k},u_{k},\lambda_{k+1})$, where $L(x_{k},u_{k},\lambda_{k+1})=0,k\neq i$. Define the Hamiltonian as follows when $k=i$
$H_{i}(x_{k},u_{k},\lambda_{k+1},\alpha_{k+1},\beta_{k})$
$=$ $L(x_{k},u_{k},\lambda_{k+1})+\alpha_{k+1}^{T}F(x_{k},u_{k})$
$\hskip 65.50012pt+\beta_{k}^{T}G(x_{k},u_{k},\lambda_{k+1}).$ (18)
Otherwise, it is given by
$H_{i}(x_{k},u_{k},\lambda_{k+1},\alpha_{k+1},\beta_{k})$
$=$ $\alpha_{k+1}^{T}F(x_{k},u_{k})+\beta_{k}^{T}G(x_{k},u_{k},\lambda_{k+1}).$ (19)
where $\alpha$ and $\beta$ are the Lagrange multiplier, $i=0,1,...,N$, and $k=0,1,...,N$. Besides, define the following FBDEs based on the Hamiltonian (18)-(19)
$\beta_{k+1}$ $=\frac{\partial H_{i}(x_{k},u_{k},\lambda_{k+1},\alpha_{k+1},\beta_{k})}{\partial\lambda_{k+1}},$ (20)
$\alpha_{k}$ $=\frac{\partial H_{i}(x_{k},u_{k},\lambda_{k+1},\alpha_{k+1},\beta_{k})}{\partial x_{k}},$ (21)
with $\beta_{0}=\textbf{0}$, $\alpha_{N+1}=\textbf{0}$.
The main result of this subsection is given below.
Theorem 1
For the OP (4), the elements of the $i$th row of the hessian matrix of the cost function $J_{N}$ with respect to $z$ are given by
$\nabla^{2}J_{N}(i)=\left[\frac{\partial H_{0}^{i}}{\partial u_{0}},\frac{\partial H_{1}^{i}}{\partial u_{1}},...,\frac{\partial H_{N}^{i}}{\partial u_{N}}\right],$ (22)
where $H^{i}_{k}=H_{i}(x_{k},u_{k},\lambda_{k+1},\alpha_{k+1},\beta_{k})$, and $x_{k},\lambda_{k+1},\alpha_{k+1}$ and $\beta_{k}$ are given by (16), (17), (20) and (21), respectively.
Proof: From the OCP (15)-(17), we formulate an optimal augmented cost function by adjoining the original cost function (15) with the FBDEs (16)-(17) using the Lagrange multiplier $\alpha_{k+1}$ and $\beta_{k}$ as
$J_{i}^{*}=\sum_{k=0}^{N}\big{\{}L(x_{k}^{*},u_{k}^{*},\lambda_{k+1}^{*})$ $+\alpha_{k+1}^{*T}[F(x_{k},u_{k})-x_{k+1}^{*}]$
$+\beta_{k}^{*T}[G(x_{k}^{*},u_{k}^{*},\lambda_{k+1}^{*})-\lambda_{k}^{*}]\big{\}}.$ (23)
Let $x_{k}$, $u_{k}$, $\lambda_{k+1}$, $x_{k+1}$, and $\lambda_{k}$ deviate from their optimal values $x_{k}^{*}$, $u_{k}^{*}$, $\lambda_{k+1}^{*}$, $x_{k+1}^{*}$, and $\lambda_{k}^{*}$ by variations $\delta x_{k}$, $\delta u_{k}$, $\delta\lambda_{k+1}$, $\delta x_{k+1}$, and $\delta\lambda_{k}$, respectively, which means that
$\xi_{k}$ $=\xi_{k}^{*}+\delta\xi_{k},$
$\zeta_{k+1}$ $=\zeta_{k+1}^{*}+\delta\zeta_{k+1}.$
where $\xi=x,u,\lambda$ and $\zeta=\lambda,x$. Based on these variations, the cost function (23) becomes
$J_{i}=\sum_{k=0}^{N-1}\big{\{}L(x_{k},u_{k},\lambda_{k+1})$ $+\alpha_{k+1}^{T}[F(x_{k},u_{k})-x_{k+1}]$
$+\beta_{k}^{T}[G(x_{k},u_{k},\lambda_{k+1})-\lambda_{k}]\big{\}}.$ (24)
It is well known that the necessary condition for optimal control is that the first variation must be equal to zero. Therefore, by applying the Taylor series expansion to (24) along with (23), it follows that
$\delta J_{i}=\sum_{k=0}^{N}\big{[}\delta x_{k}^{T}\frac{\partial V^{i}_{k}}{\partial x_{k}^{*}}+\delta u_{k}^{T}\frac{\partial V^{i}_{k}}{\partial u_{k}^{*}}+\delta\lambda_{k+1}^{T}\frac{\partial V^{i}_{k}}{\partial\lambda_{k+1}^{*}}$
$+\delta x_{k+1}^{T}\frac{\partial V^{i}_{k}}{\partial x_{k+1}^{*}}+\delta\lambda_{k}^{T}\frac{\partial V^{i}_{k}}{\partial\lambda_{k}^{*}}\big{]},$ (25)
where
$V_{k}^{i}$ $=V_{i}(x_{k}^{*},u_{k}^{*},\lambda_{k+1}^{*},\alpha_{k+1}^{*},x_{k+1}^{*},\beta_{k}^{*},\lambda_{k}^{*})$
$=L(x_{k}^{*},u_{k}^{*},\lambda_{k+1}^{*})+\alpha_{k+1}^{*T}[F(x_{k},u_{k})-x_{k+1}^{*}]$
$\hskip 81.00018pt+\beta_{k}^{*T}[G(x_{k}^{*},u_{k}^{*},\lambda_{k+1}^{*})-\lambda_{k}^{*}].$
To express the variation $\delta x_{k+1}$ in terms of $\delta x_{k}$, examine the fourth expression in (25) below.
$\sum_{k=0}^{N}\delta x_{k+1}^{T}\frac{\partial V^{i}_{k}}{\partial x_{k+1}^{*}}=\sum_{k=1}^{N+1}\delta x_{k}^{T}\frac{\partial V^{i}_{k-1}}{\partial x_{k}^{*}}$
$=$ $\sum_{k=0}^{N}\delta x_{k}^{T}\frac{\partial V^{i}_{k-1}}{\partial x_{k}^{*}}+(\delta x_{N+1}^{T}\frac{\partial V^{i}_{N}}{\partial x_{N+1}^{*}}-\delta x_{0}^{T}\frac{\partial V^{i}_{-1}}{\partial x_{0}^{*}}).$ (26)
Similarly, consider the fifth expression in (25) below.
$\sum_{k=0}^{N}\delta\lambda_{k}^{T}\frac{\partial V^{i}_{k}}{\partial\lambda_{k}^{*}}=\sum_{k=-1}^{N-1}\delta\lambda_{k+1}^{T}\frac{\partial V^{i}_{k+1}}{\partial\lambda_{k+1}^{*}}$
$=$ $\sum_{k=0}^{N}\delta\lambda_{k+1}^{T}\frac{\partial V^{i}_{k+1}}{\partial\lambda_{k+1}^{*}}+(\delta\lambda_{0}^{T}\frac{\partial V^{i}_{0}}{\partial\lambda_{0}^{*}}-\delta\lambda_{N+1}^{T}\frac{\partial V^{i}_{N+1}}{\partial\lambda_{N+1}^{*}}).$ (27)
Substituting the expressions (26)-(27) into (25), the first variation $\delta J_{i}$ can be reformulate as
$\delta J_{i}=\sum_{k=0}^{N}$ $\big{[}\delta x_{k}^{T}(\frac{\partial V^{i}_{k}}{\partial x_{k}^{*}}+\frac{\partial V^{i}_{k-1}}{\partial x_{k}^{*}})+\delta\lambda_{k+1}^{T}(\frac{\partial V^{i}_{k}}{\partial\lambda_{k+1}^{*}}+\frac{\partial V^{i}_{k+1}}{\partial\lambda_{k+1}^{*}})$
$+\delta u_{k}^{T}\frac{\partial V^{i}_{k}}{\partial u_{k}^{*}}\big{]}+(\delta x_{N+1}^{T}\frac{\partial V^{i}_{N}}{\partial x_{N+1}^{*}}-\delta x_{0}^{T}\frac{\partial V^{i}_{-1}}{\partial x_{0}^{*}})$
$+(\delta\lambda_{0}^{T}\frac{\partial V^{i}_{0}}{\partial\lambda_{0}^{*}}-\delta\lambda_{N+1}^{T}\frac{\partial V^{i}_{N+1}}{\partial\lambda_{N+1}^{*}}).$ (28)
Combining the Hamiltonian (18) and (19), we have
$V_{k}^{i}=H_{i}(x_{k}^{*},u_{k}^{*},\lambda_{k+1}^{*},\alpha_{k+1}^{*},\beta_{k}^{*})-\alpha_{k+1}^{*T}x^{*}_{k+1}-\beta_{k}^{*T}\lambda_{k}.$
Taking into account the arbitrary nature of the variations in (28), we can obtain
$\frac{\partial H_{i}(x_{k}^{*},u_{k}^{*},\lambda_{k+1}^{*},\alpha_{k+1}^{*},\beta_{k}^{*})}{\partial u_{k}^{*}}=0,$ (29)
where $x_{k}^{*},\lambda_{k+1}^{*},\alpha_{k+1}^{*}$ and $\beta_{k}^{*}$ satisfy (16), (17), (20) and (21), respectively. When the given $u_{k}$ does note meet the equilibrium condition (29), the left side of equation (29) represents the $k$th element of the gradient, i.e., the (22) represents the $i$th row of the hessian matrix of the cost function $J_{N}$. The proof is complete.
Building on the three steps of the gradient computation, the implementation steps of Theorem 1 are summarized as follows:
Step 1: For the given $x_{k}$, $u_{k}$, and $\lambda_{k}$, compute the forward difference equation (20) from $k=0$ to $k=N-1$ to obtain the $\beta_{k},k=0,1,...,N$ using the given initial value $\beta_{0}$.
Step 2: Compute the backward difference equation (21) from $k=N$ to $k=1$ to obtain the $\alpha_{k},k=1,...,N+1$ using the given terminal value $\alpha_{N+1}$.
Step 3: Compute the elements of the $i$th row of the hessian matrix of $J_{N}$ according to (22).
Step 4: Repeat steps 1 to 3 until all elements of the hessian matrix are computed.
Algorithm 1 Numerical algorithm for solving the OCP (1)-(2).
1:Initialize $x_{t_{0}}$, $z_{0}=[u_{t_{0}}^{T},...,u_{t_{N_{p}}}^{T}]^{T}$;
2:Optimal controller $u_{t_{0}}$, $t_{0}=0,...,N$;
3:for $k=0$ to $N$ do
4:     for $i=0$ to #$\#step$ do
5:         Calculate $\nabla J=\nabla J_{N}(x_{t_{0}},z_{i})$;
6:         Calculate $\nabla^{2}J=\nabla^{2}J_{N}(x_{t_{0}},z_{i})$;
7:         $g_{0}(z_{i})=(R+\nabla^{2}J)^{-1}\nabla J$;
8:         for $j=1$ to $i$ do
9:              $g_{j}(z_{i})=(R+\nabla^{2}J)^{-1}[\nabla J+Rg_{j-1}(z_{i})]$;
10:         end for
11:         $z_{i+1}=z_{i}-g_{i}(z_{i})$;
12:     end for
13:     Extract $u_{t_{0}}$ from $z_{i+1}$;
14:     Update initial state $x_{t_{0}}$;
15:end for
III-D Implementation
In long-horizon OCP, performance is often compromised by numerical errors and model inaccuracies. MPC mitigates these issues by solving a sequence of local optimal control problems online. Instead of optimizing over the entire time horizon in one step, MPC focuses on a finite prediction horizon $k=t_{0},...,t_{N_{p}}$ at each iteration, updating the control inputs based on the current system state.
In practical implementation, solving the OCP (1)-(2) at each sampling time is necessary within the MPC framework. During the optimization process, firstly, compute the gradient of the cost function with respect to the control variables using Lemma 1. Secondly, compute the hessian matrix of the cost function with respect to the control variables as described in Theorem 1. Finally, iteratively update the optimization algorithm by applying formulas (9) through (11). Repeat this process until the optimization converges. The details are summarized in Algorithm 1.
IV Conclusion
This paper presents a method that combines nonlinear optimization with the Pontryagin’s maximum principle to solve discrete-time nonlinear OCP. The optimization objective is redefined from a new optimal control perspective, and a fast-converging iterative algorithm is designed, developing a novel approach for solving nonlinear OCP using optimal control theory. Additionally, explicit formulas based on FBDEs are derived to enhance the computational efficiency of evaluating gradient and second-order derivative for complex cost function.